# sources.bib

### v1

@@ 3. Group-graded twisted Calabi–Yau algebras
Just as [zgrad] understands the first Weyl algebra as its “prototypical” example, so too do we take this algebra as our motivating example of a group-graded Calabi–Yau algebra.
@@ 3.1. Definition and Compatibility with Bimodule Structure
Let $A$ be a $G$-graded $k$-algebra. We say that $A$ is $G$-graded twisted Calabi–Yau of dimension $d$ if $A$ is graded homologically smooth and there exists $U$ such that the following holds as a graded isomorphism of $(A,A)$-bimodules:
$\operatorname{\underline{EXT}}_{A^{e}}^{i}(A,A^{e})\cong\begin{cases}0,\ i\neq d\\ U,\ i=d\end{cases}$
A few comments are in order. Consider ${}^{1}A^{\nu}$, the graded $(A,A)$-bimodule $A$, with normal left action, but right action twisted by $\nu$, a graded algebra automorphism of $A$. Should Definition 3.1 hold with $U\cong{{}^{1}}A^{\nu}(\sigma)$ for some $\sigma\in G$, then we say that $A$ has the graded Nakayama automorphism and is of Gorenstein index $\sigma$. Should it further hold with $\nu$ the identity map, then we say that $A$ is simply $G$-graded (untwisted) Calabi–Yau.
@@
Example 1.
[Weyl Algebras with Standard $\mathbb{Z}$-grading, $\mathbb{Z}^{n}$-grading] We consider the $n$th Weyl algebra, $A_{n}$, defined as follows:
@@
$0\to A_{n}\otimes\bigwedge^{2n}V\otimes A_{n}\to A_{n}\otimes\bigwedge^{2n-1}V\otimes A_{n}\to\dots\to A_{n}\otimes A_{n}\twoheadrightarrow{}A_{n}\to 0$
$f\otimes(v_{1}\wedge\dots\wedge v_{i})\otimes g\mapsto\sum_{j=1}^{i}$ $(-1)^{j-1}(fv_{j})\otimes(v_{1}\wedge\dots\hat{v}_{j}\dots\wedge v_{i})\otimes g$
$+$ $(-1)^{j}f\otimes(v_{1}\wedge\dots\hat{v}_{j}\dots\wedge v_{i})\otimes(v_{j}g)$
@@
We need now argue that the differentials are morphisms in the graded module category, which amounts to showing that they are degree-preserving. Consider $f,g\in A_{n}$ homogeneous of degrees $m_{1},m_{2}$, respectively, and $v=v_{1}\wedge\dots\wedge v_{i}\in\bigwedge^{i}V$ such that each $v_{k}$ is homogeneous of degree $l_{k}$, $1\leq k\leq i$. Then $f\otimes(v_{1}\wedge\dots\wedge v_{n})\otimes g$ is of degree $m_{1}+\sum_{j=1}^{i}l_{j}+m_{2}=m_{1}+|v|+m_{2}$ in $A_{n}\otimes\bigwedge^{i}V\otimes A_{n}$. Now, considering the image of the differential above, notice that each term in the summand preserves the degree. Indeed, for any $1\leq j\leq i$, we see
$(-1)^{j-1}\underbrace{(fv_{j})}_{\deg m_{1}+l_{j}}$ $\otimes\underbrace{(v_{1}\wedge\dots\hat{v}_{j}\dots\wedge v_{i})}_{\deg|v|-l_{j}}\otimes\underbrace{g}_{\deg m_{2}}$
$+(-1)^{j}\underbrace{f}_{\deg m_{2}}$ $\otimes\underbrace{(v_{1}\wedge\dots\hat{v}_{j}\dots\wedge v_{i})}_{\deg|v|-l_{j}}\otimes\underbrace{(v_{j}g)}_{|v|+m_{2}}$
has total degree $m_{1}+|v|+m_{2}$ in $A_{n}\otimes\bigwedge^{i-1}V\otimes A_{n}$. Hence, the above resolutions provides a projective resolution of $A_{n}$ in the category of $\mathbb{Z}$-graded $A^{e}$-modules. It immediately follows that $\operatorname{\underline{EXT}}^{2n}_{A^{e}}(A,A^{e})\cong A_{n}$ in this category. We conclude that $A_{n}$ is $\mathbb{Z}$-graded Calabi–Yau of Gorenstein index 00 for all $n$.
@@
Example 2 (Generalized Weyl Algebras).
First introduced by Bavula in [bavula1992generalized] and later generalized in [bavula1996tensor], the generalized Weyl algebras (GWAs) of degree $n$ over a ring $D$ are defined as follows. Let $\sigma=(\sigma_{1},\dots,\sigma_{n})$ a tuple of $n$ commuting automorphisms of $D$, and $a=(a_{1},\dots,a_{n})$ a tuple of $n$ central elements of $D$ such that $\sigma_{i}(a_{j})=a_{j}$ for all $i\neq j$. Then the degree $n$ GWA over $D$ is the ring $A=D(\sigma,a)$ generated by $D$ and $2n$ indeterminates $x_{1},\dots,x_{n},y_{1},\dots,y_{n}$ which are subject to the relations
@@
$[x_{i},x_{j}]$ $=[y_{i},y_{j}]=[x_{i},y_{j}]=0\quad\forall i\neq j,$
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the commutator. Any degree $n$ GWA further possesses a $\mathbb{Z}^{n}$-grading by setting $|x_{i}|=\mathbf{e_{i}}$ and $|y_{i}|=-\mathbf{e_{i}}$ (c.f.[bavula1992generalized]).
As in [Liu-HomSmoothGWAI], consider $D=k[z]$, $\sigma=\sigma_{1}=\lambda z+\eta$ for $\lambda\neq 0,\eta\in k$, and $a=a_{1}$ a polynomial in $k[z]$. Theorem 1.1 of [Liu-HomSmoothGWAI] tells us that the resulting degree $1$ GWA $A=D(\sigma,a)$ is homologically smooth, and in particular twisted Calabi–Yau of dimension $2$, if and only if $a$ has no multiple roots. As in our above discussion, $A$ also has a $\mathbb{Z}$-grading, and is hence $\mathbb{Z}$-graded twisted Calabi–Yau of dimension $2$ by Theorem 3.2. Theorem 1.1 of [LIU2018228] also gives conditions under which a GWA over $D=k[z_{1},z_{2}]$ is twisted Calabi–Yau, which again by our previous discussions, yields examples of $\mathbb{Z}^{2}$-graded twisted Calabi–Yau algebras. It would be interesting to see more examples of Calabi–Yau algebras arising from GWAs over more general rings $D$. As noted above, all GWAs possess a $\mathbb{Z}^{n}$-grading, so this would in turn produce more examples of group-graded twisted Calabi–Yau algebras.
Example 3 (Universal Enveloping Algebras of Finite-Dimensional Lie Algebras).
@@ 4. Localizations of twisted Calabi–Yau algebras
It is well-known that $k[t]$ provides an example of a $1$-dimensional $\mathbb{N}$-graded Calabi–Yau algebra, and we might observe that localizing at $t$ turns this connected, $\mathbb{N}$-graded algebra into the connected, strongly $\mathbb{Z}$-graded algebra $k[t,t^{-1}]$. This prompts us to wonder if $k[t,t^{-1}]$ is Calabi–Yau, and more generally ask if localization always preserves the twisted Calabi–Yau property. If this were to be true, we would find a rich source of $\mathbb{Z}$-graded Calabi–Yau algebras as localizations of Artin–Schelter regular algebras.
@@
Let $A,B$ $k$-algebras such that $S_{1}\subseteq A$ is a left (resp. right) denominator subset, and such that $S_{2}\subseteq B$ is also a left (resp. right) denominator subset. Then,
@@
Let $S_{1},S_{2},$ and $T$ as above. We will show that $T$ is a left denominator set, and leave the proof on the right as an analogous argument. It is clear that $T$ is a multiplicative set, so it suffices to show that $T$ is left permutable and left reversible.
Towards left permutability, fix $s_{1}\otimes s_{2}\in T$ and $a\otimes b\in A\otimes B$. Recall that since $S_{1}\subseteq A$ and $S_{2}\subseteq B$ are left denominator sets, for all $a\in A,b\in B,s_{1}\in S_{1},$ and $s_{2}\in S_{2}$ we have
[…]
from which one can show
$T(a\otimes b)\cap(A\otimes B)(s_{1}\otimes s_{2})$ $=(S_{1}a\otimes S_{2}b)\cap(As_{1}\otimes Bs_{2})\neq\emptyset\text{,}$
and safely conclude that $T$ is left permutable. Let us now turn our attention to showing that $T$ is left reversible. Towards this end, fix $a\otimes b\in A\otimes B$, $s_{1}\otimes s_{2}\in T$, and suppose that
$(a\otimes b)(s_{1}\otimes s_{2})=as_{1}\otimes bs_{2}=0\text{.}$
Recall that since $A\otimes B$ is a tensor product over $k$ of two $k$-vector spaces, we must have that the above implies either $as_{1}=0$ or $bs_{2}=0$. If $as_{1}=0$, then left reversibility of $S_{1}\subseteq A$ gives the existence of $\hat{s}_{1}\in S_{1}$ such that $\hat{s}_{1}a=0$. Consider $\hat{s}_{1}\otimes 1\in T$. Then,
$(\hat{s}_{1}\otimes 1)(a\otimes b)$ $=(\hat{s}_{1}a\otimes b)$
$=0\otimes b=0\text{,}$
thus yielding left reversibility, in this case. The result in the case where $bs_{2}=0$ follows similarly from left reversibility of $S_{2}$. We see that $T$ indeed remains left reversible, and may now conclude that $T$ provides the desired left denominator subset of $A\otimes B$. ∎
We are most interested in the case where we have $A$ a $k$-algebra viewed as a left $A^{e}$-module. If $S$ is a left (resp. right) denominator subset of $A$, then we can consider the localization of $A$ as a left $A$-module, $S^{-1}A$ (resp. as a right $A$-module $AS^{-1}$). To make sense of the localization of $A$ viewed as a left $A^{e}$-module, we want to consider $T=S\otimes S^{\operatorname{op}}$ for $S\subseteq A$ a left and right denominator set, as in Lemma 4.1.
Lemma 4.2.
Let $A$ and $B$ be $k$-algebras. Let $S_{1}\subseteq A$ a left and right denominator subset, $S_{2}\subseteq B$ a left and right denominator subset, and define $T\coloneqq S_{1}\otimes S_{2}$ as in the Lemma 4.1. Then $T^{-1}(A\otimes B)\cong S_{1}^{-1}A\otimes S_{2}^{-1}B$ as $k$-algebras.
Proof.
We will show that $S_{1}^{-1}A\otimes S_{2}^{-1}B$ satisfies the universal property of $A\otimes B$ localized at $T$. Towards that end, consider $\psi:A\otimes B\to R$ a $T$-inverting morphism of $k$-algebras and $i_{A,B}:A\ (\text{resp. }B)\to A\otimes B$ the canonical injections. By the universal property of the tensor product of algebras,
($\heartsuit$) $\operatorname{Hom}(A\otimes B,R)\cong\{(f,g)\in\operatorname{Hom}(A,R)\times\operatorname{Hom}(B,R):[f(A),g(B)]=0\},$
where $[f(A),g(B)]=0$ means that for all $a\in A$ and $b\in B$ we have $f(a)g(b)-g(b)f(a)=0$, whence $\psi$ corresponds to $(\psi\circ i_{A},\psi\circ i_{B})$. Since $i_{A}(S_{1})=S_{1}\otimes 1\subseteq T$, we observe that $\psi\circ i_{A}:A\to R$ is an $S_{1}$-inverting $k$-algebra homomorphism, and similarly we see $\psi\circ i_{B}:B\to R$ an $S_{2}$-inverting $k$-algebra homomorphism.
Now, by the universal property of localization of $A$, we know there exists a unique $k$-algebra homomorphism $\varphi_{1}:S_{1}^{-1}A\to R$, making the squares in the diagram below commute (an analogous argument shows the same holds true for $\varphi_{2}:S_{2}^{-1}B\to R$).
${S_{1}^{-1}A\otimes S_{2}^{-1}B}$${A}$${A\otimes B}$${B}$${S_{1}^{-1}A}$${R}$${S_{2}^{-1}B}$$\scriptstyle{\hat{\varphi}}$$\scriptstyle{i_{A}}$$\scriptstyle{\psi}$$\scriptstyle{i_{b}}$$\scriptstyle{\varphi_{1}}$$\scriptstyle{\varphi_{2}}$
Again by the universal property of tensor product, we know that morphisms from $S_{1}^{-1}A\otimes S_{2}^{-1}B$ are determined by pairs of morphisms from $S_{1}^{-1}A\to R$ and $S_{2}^{-1}B\to R$ whose images centralize each other. We already have $\varphi_{1},\varphi_{2}$ giving such maps, so it suffices to check that they centralize each other. This is readily seen by recalling that $\varphi_{1}(S_{1}^{-1}A)=S_{1}^{-1}A\otimes 1$ and $\varphi_{2}(S_{2}^{-1}B)=S_{2}^{-1}B\otimes 1$, and observing that these images centralize each other.
Recalling the characterization $\heartsuit$4 above, $\hat{\varphi}$ is uniquely determined by $(\varphi_{1},\varphi_{2})$, and so it provides an honest $k$-algebra homomorphism from $S_{1}^{-1}A\otimes S_{2}^{-1}B$ to $R$. It is easy to see that $\hat{\varphi}$ is the unique $k$-algebra homomorphism giving commutativity of the desired subdiagram. We conclude that $S_{1}^{-1}A\otimes S_{2}^{-1}B$ satisfies the same universal property of $T^{-1}(A\otimes B)$, so the desired isomorphism holds. ∎
Lemma 4.3.
Let $A$ be an algebra over a field $k$. If $A$ is homologically smooth, then $S^{-1}A$ is homologically smooth, where $S$ is both a left and right denominator set of $A$.
Proof.
@@
Let $A$ as above. Since $A$ is homologically smooth, there exists a finite length projective resolution $P_{\bullet}\to A\to 0$ of finitely generated projective $A^{e}$-modules. Let $S\subseteq A$ a left and right denominator subset. It is clear that $S^{\operatorname{op}}\subseteq A^{\operatorname{op}}$ is a left and right denominator subset from our observations above, so we may define the left and right denominator subset $\mathfrak{S}:=S\otimes S^{\operatorname{op}}\subseteq A\otimes A^{\operatorname{op}}=A^{e}$ as in Lemma 4.1 above. Let us now localize $P_{\bullet}\to A\to 0$ at $\mathfrak{S}$. Since localization is exact, $\mathfrak{S}^{-1}P_{\bullet}\to\mathfrak{S}^{-1}A\to 0$ remains an exact sequence of $\mathfrak{S}^{-1}A^{e}$-modules. It is well known that localizations of finitely generated projective modules remain finitely generated projectives modules over the localized ring; hence, $\mathfrak{S}^{-1}P_{\bullet}\to\mathfrak{S}^{-1}A\to 0$ is a projective resolution of finitely generated $\mathfrak{S}^{-1}A^{e}$-modules.
@@
$\cong\left(S^{-1}A\otimes_{k}(S^{-1}A)^{\operatorname{op}}\right)\otimes_{A^{e}}A$
@@
[Localization and Hochschild (Co)homology, cf. [brylinski1989central]] Let $A$ be a $k$-algebra which is finitely presented as a left $A^{e}$-module and $M$ an $A^{e}$-module. Suppose further that $S\subseteq A$ is a left and right denominator set so that we may define $\mathfrak{S}:=S\otimes S^{\operatorname{op}}\subseteq A^{e}$ as in Lemma 4.3. Then, the following hold as isomorphisms of $\mathfrak{S}^{-1}A^{e}$-modules:
@@
$\operatorname{Ext}^{i}_{\mathfrak{S}^{-1}A^{e}}(\mathfrak{S}^{-1}A,\mathfrak{S}^{-1}M)$ $\hphantom{:}=:\text{HH}^{i}(S^{-1}A,\mathfrak{S}^{-1}M)$
@@
Now, it is well-known that because $\mathfrak{S}^{-1}A^{e}$ is flat over $A^{e}$ and $A$ finitely presented as an $A^{e}$-module, then
@@
Let $0\to P_{n}\to\dots\to P_{0}\to A$ be a resolution of finitely generated and projective $A^{e}$-modules that we get from $A$ being homologically smooth. In particular, we have the short exact sequence $0\to\ker(d_{0})\to P_{0}\to A\to 0$. Notice that $P_{0}$ is finitely presented as a left $A^{e}$-module, and $\ker(d_{0})=\operatorname{im}(d_{1})$. Since the homomorphic image of a finitely generated module is finitely generated, we have that $\ker(d_{0})$ is finitely generated. We conclude that $A$ is finitely presented as a left $A^{e}$-module, as desired. ∎
@@
  1. (1),(3),(4)
    Since $U$ and $V$ are invertible $(A,A)$-bimodules, $U\cong\operatorname{Hom}_{A}(V,A)$ and $V\cong\operatorname{Hom}(U,A)$ hold as isomorphisms of $(A,A)$-bimodules. Moving $AS^{-1}$ inside the parentheses is justified by [eisenbud, Proposition 2.10], the proof of which holds in the noncommutative setting. Recall that $V$ and $U$ are invertible bimodules, and are thus finitely generated and projective $A$ modules on the left and right. In particular, they are finitely presented as left and right $A$-modules.
  2. (2)
    Notice that $V$ is finitely generated and projective as a left $A$-modules, so $S^{-1}V$ remains so as a left $S^{-1}A$-module. Now apply [sacred3, Section 2, Exercise 20].
@@
Let $A$ a $k$-algebra and $S$ a left and right denominator subset of $A$. If $A$ is twisted Calabi–Yau of dimension $d$, then $S^{-1}A$ is twisted Calabi–Yau of dimension $d$.
@@
(5) $\mathfrak{S}^{-1}U\otimes_{S^{-1}A}\mathfrak{S}^{-1}V$ $\cong S^{-1}U\otimes_{S^{-1}A}S^{-1}V$
(6) $\cong S^{-1}(U\otimes_{A}V)$
(7) $\cong S^{-1}A.$
Recall that (5) is Lemma 4.6 applied to the same argument used in Lemma 4.3 to argue that $\mathfrak{S}^{-1}A\cong S^{-1}A$ as $(S^{-1}A)^{e}$-modules.

### v2

@@ 3.1. Definition and Compatibility with Bimodule Structure
Let $A$ be a $G$-graded $k$-algebra. We say that $A$ is $G$-graded twisted Calabi–Yau of dimension $d$ if $A$ is graded homologically smooth and there exists a $G$-graded invertible $(A,A)$-bimodule $U$ such that the following holds as a graded isomorphism of $(A,A)$-bimodules:
$\operatorname{\underline{EXT}}_{A^{e}}^{i}(A,A^{e})\cong\begin{cases}0,\ i\neq d;\\ U,\ i=d.\end{cases}$
A few comments are in order. Consider ${}^{1}A^{\nu}$, the graded $(A,A)$-bimodule $A$, with normal left action, but right action twisted by $\nu$, a graded algebra automorphism of $A$. Should Definition 3.1 hold with $U\cong{{}^{1}}A^{\nu}(\sigma)$ for some $\sigma\in G$, then we say that $\nu$ is a graded Nakayama automorphism of $A$ and is of Gorenstein index $\sigma$. Should it further hold with $\nu$ the identity map, then we say that $A$ is simply $G$-graded (untwisted) Calabi–Yau.
@@
Example 3.3.
[Weyl Algebras with Standard $\mathbb{Z}$-grading, $\mathbb{Z}^{n}$-grading] Let $k$ be a field of characteristic zero. We consider the $n$th Weyl algebra, $A_{n}$, defined as follows:
@@
$0\to A_{n}\otimes\bigwedge^{2n}V\otimes A_{n}\to A_{n}\otimes\bigwedge^{2n-1}V\otimes A_{n}\to\cdots\to A_{n}\otimes A_{n}\twoheadrightarrow{}A_{n}\to 0,$
$f\otimes(v_{1}\wedge\cdots\wedge v_{i})\otimes g\mapsto\sum_{j=1}^{i}$ $(-1)^{j-1}(fv_{j})\otimes(v_{1}\wedge\cdots\hat{v}_{j}\cdots\wedge v_{i})\otimes g$
$+$ $(-1)^{j}f\otimes(v_{1}\wedge\cdots\hat{v}_{j}\cdots\wedge v_{i})\otimes(v_{j}g).$
@@
We need now argue that the differentials are morphisms in the graded module category, which amounts to showing that they are degree-preserving. Consider $f,g\in A_{n}$ homogeneous of degrees $m_{1},m_{2}$, respectively, and $v=v_{1}\wedge\dots\wedge v_{i}\in\bigwedge^{i}V$ such that each $v_{k}$ is homogeneous of degree $l_{k}$, $1\leq k\leq i$. Then $f\otimes(v_{1}\wedge\dots\wedge v_{i})\otimes g$ is of degree $m_{1}+\sum_{j=1}^{i}l_{j}+m_{2}=m_{1}+|v|+m_{2}$ in $A_{n}\otimes\bigwedge^{i}V\otimes A_{n}$. Now, considering the image of the differential above, notice that each term in the summand preserves the degree. Indeed, for any $1\leq j\leq i$, we see
$(-1)^{j-1}\underbrace{(fv_{j})}_{\deg m_{1}+l_{j}}$ $\otimes\underbrace{(v_{1}\wedge\cdots\hat{v}_{j}\cdots\wedge v_{i})}_{\deg|v|-l_{j}}\otimes\underbrace{g}_{\deg m_{2}}$
$+(-1)^{j}\underbrace{f}_{\deg m_{1}}$ $\otimes\underbrace{(v_{1}\wedge\cdots\hat{v}_{j}\cdots\wedge v_{i})}_{\deg|v|-l_{j}}\otimes\underbrace{(v_{j}g)}_{\deg|v|+m_{2}}$
has total degree $m_{1}+|v|+m_{2}$ in $A_{n}\otimes\bigwedge^{i-1}V\otimes A_{n}$. Hence, the above resolution provides a projective resolution of $A_{n}$ in the category of $\mathbb{Z}$-graded $A^{e}$-modules. It immediately follows that $\operatorname{\underline{EXT}}^{2n}_{A^{e}}(A,A^{e})\cong A_{n}$ in this category. We conclude that $A_{n}$ is $\mathbb{Z}$-graded Calabi–Yau of Gorenstein index 00 for all $n$.
@@
Example 3.4 (Generalized Weyl Algebras).
First introduced by Bavula in [bavula1992generalized] and later generalized in [bavula1996tensor], the generalized Weyl algebras (GWAs) of degree $n$ over a ring $D$ are defined as follows. Let $\sigma=(\sigma_{1},\dots,\sigma_{n})$ be a tuple of $n$ commuting automorphisms of $D$, and let $a=(a_{1},\dots,a_{n})$ be a tuple of $n$ central elements of $D$ such that $\sigma_{i}(a_{j})=a_{j}$ for all $i\neq j$. Then the degree $n$ GWA over $D$ is the ring $A=D(\sigma,a)$ generated by $D$ and $2n$ indeterminates $x_{1},\dots,x_{n},y_{1},\dots,y_{n}$ which are subject to the relations
@@
$[x_{i},x_{j}]$ $=[y_{i},y_{j}]=[x_{i},y_{j}]=0\quad\forall i\neq j.$
Any degree $n$ GWA further possesses a $\mathbb{Z}^{n}$-grading by setting $|x_{i}|=\mathbf{e_{i}}$ and $|y_{i}|=-\mathbf{e_{i}}$ (cf. [bavula1992generalized]).
As in [Liu-HomSmoothGWAI], consider $D=k[z]$, $\sigma=\sigma_{1}=\lambda z+\eta$ for $\lambda\neq 0,\eta\in k$, and $a=a_{1}$ a polynomial in $k[z]$. Theorem 1.1 of [Liu-HomSmoothGWAI] tells us that the resulting degree $1$ GWA $A=D(\sigma,a)$ is homologically smooth, and in particular twisted Calabi–Yau of dimension $2$, if and only if $a$ has no multiple roots. As in our above discussion, $A$ also has a $\mathbb{Z}$-grading, and is hence $\mathbb{Z}$-graded twisted Calabi–Yau of dimension $2$ by Theorem 3.2. Further, [LIU2018228, Theorem 1.1] gives conditions under which a GWA over $D=k[z_{1},z_{2}]$ is twisted Calabi–Yau, which again by our previous discussions, yields examples of $\mathbb{Z}^{2}$-graded twisted Calabi–Yau algebras. It would be interesting to see more examples of Calabi–Yau algebras arising from GWAs over more general rings $D$. As noted above, all GWAs possess a $\mathbb{Z}^{n}$-grading, so this would in turn produce more examples of group-graded twisted Calabi–Yau algebras.
Example 3.5 (Universal Enveloping Algebras of Finite-Dimensional Lie Algebras).
@@ 4. Localizations of twisted Calabi–Yau algebras
It is well-known that $k[t]$ provides an example of a $1$-dimensional $\mathbb{N}$-graded Calabi–Yau algebra, and we can observe that localizing at $t$ turns this connected, $\mathbb{N}$-graded algebra into the connected, strongly $\mathbb{Z}$-graded algebra $k[t,t^{-1}]$. We will show that localization at left and right denominator subsets preserves the twisted Calabi–Yau property, which in turn will give us a new way of finding group-graded twisted Calabi–Yau algebras.
Indeed, closely related to the study of Calabi–Yau algebras is that of Artin–Schelter regular algebras, a definition of which can be found in [rogalski2023artinschelter], for example. For our purposes, we will take the definition of an Artin–Schelter regular algebra to be a connected $\mathbb{N}$-graded twisted Calabi–Yau algebra, as they were proven to be equivalent in [Reyes_2014, Lemma 1.2]. As such, we will find a rich source of $\mathbb{Z}$-graded twisted Calabi–Yau algebras arising as localizations of Artin–Schelter regular algebras.
We start by stating two well-known technical lemmas, the proofs of which we omit. Recall that a multiplicative subset $S$ of a ring $R$ is said to be a left (resp. right) denominator subset if it is left (resp. right) permutable and left (resp. right) reversible. That is,
  • for all $a\in R$ and $s\in S$, $Sa\cap Rs\neq\emptyset$ (resp. $aS\cap sR\neq\emptyset$),
  • for all $a\in R$, if $as^{\prime}=0$ for some $s^{\prime}\in S$, then $sa=0$ for some $s\in S$ (resp. if $s^{\prime}a=0$ for some $s^{\prime}\in S$, then $as=0$ for some $s\in S$).
@@
Let $A$ and $B$ be $k$-algebras such that $S_{1}\subseteq A$ is a left (resp. right) denominator subset, and such that $S_{2}\subseteq B$ is also a left (resp. right) denominator subset. Then,
@@
We are most interested in the case where we have $A$ a $k$-algebra viewed as a left $A^{e}$-module. If $S$ is a left (resp. right) denominator subset of $A$, then we can consider localization of $A$ as a left $A$-module, $S^{-1}A$ (resp. as a right $A$-module $AS^{-1}$). To make sense of the localization of $A$ viewed as a left $A^{e}$-module, we want to consider $T=S\otimes S^{\operatorname{op}}$ for $S\subseteq A$ a left and right denominator set, as in Lemma 4.1.
Lemma 4.2.
Let $A$ and $B$ be $k$-algebras. Let $S_{1}\subseteq A$ be a left and right denominator subset, let $S_{2}\subseteq B$ be a left and right denominator subset, and define $T\coloneqq S_{1}\otimes S_{2}$ as in the Lemma 4.1. Then $T^{-1}(A\otimes B)\cong S_{1}^{-1}A\otimes S_{2}^{-1}B$ as $k$-algebras.
Lemma 4.3.
Let $A$ be a $k$-algebra. If $A$ is homologically smooth, then $S^{-1}A$ is homologically smooth, where $S$ is both a left and right denominator set of $A$.
@@
Let $A$ be as above. Since $A$ is homologically smooth, there exists a finite length projective resolution $P_{\bullet}\to A\to 0$ of finitely generated projective $A^{e}$-modules. Let $S\subseteq A$ be a left and right denominator subset. It is clear that $S^{\operatorname{op}}\subseteq A^{\operatorname{op}}$ is a left and right denominator subset from our observations above, so we may define the left and right denominator subset $\mathfrak{S}:=S\otimes S^{\operatorname{op}}\subseteq A\otimes A^{\operatorname{op}}=A^{e}$ as in Lemma 4.1. Let us now localize $P_{\bullet}\to A\to 0$ at $\mathfrak{S}$. Since localization is exact, $\mathfrak{S}^{-1}P_{\bullet}\to\mathfrak{S}^{-1}A\to 0$ remains an exact sequence of $\mathfrak{S}^{-1}A^{e}$-modules. It is well known that localizations of finitely generated projective modules remain finitely generated projective modules over the localized ring; hence, $\mathfrak{S}^{-1}P_{\bullet}\to\mathfrak{S}^{-1}A\to 0$ is a projective resolution of finitely generated $\mathfrak{S}^{-1}A^{e}$-modules.
@@
$\cong\left(S^{-1}A\otimes(S^{-1}A)^{\operatorname{op}}\right)\otimes_{A^{e}}A$
@@
[Localization and Hochschild (Co)homology] Let $A$ be a $k$-algebra which is finitely presented as a left $A^{e}$-module, and let $M$ be an $A^{e}$-module. Suppose further that $S\subseteq A$ is a left and right denominator set so that we may define $\mathfrak{S}:=S\otimes S^{\operatorname{op}}\subseteq A^{e}$ as in Lemma 4.3. Then, the following hold as isomorphisms of $\mathfrak{S}^{-1}A^{e}$-modules:
@@
$\operatorname{Ext}^{i}_{\mathfrak{S}^{-1}A^{e}}(\mathfrak{S}^{-1}A,\mathfrak{S}^{-1}M)$ $\hphantom{:}=:\text{HH}^{i}(S^{-1}A,\mathfrak{S}^{-1}M).$
@@
Now, it is well-known that because $\mathfrak{S}^{-1}A^{e}$ is flat over $A^{e}$ and $A$ is finitely presented as an $A^{e}$-module, then
@@
Since $A$ is homologically smooth, there is a resolution $0\to P_{n}\to\dots\to P_{1}\xrightarrow{d_{1}}P_{0}\xrightarrow{d_{0}}A\to 0$ of finitely generated and projective $A^{e}$-modules. In particular, we have the short exact sequence $0\to\ker(d_{0})\to P_{0}\to A\to 0$. Notice that $P_{0}$ is finitely presented as a left $A^{e}$-module, and $\ker(d_{0})=\operatorname{im}(d_{1})$. Since the homomorphic image of a finitely generated module is finitely generated, we have that $\ker(d_{0})$ is finitely generated. We conclude that $A$ is finitely presented as a left $A^{e}$-module, as desired. ∎
@@
(1,3,4) Since $U$ and $V$ are invertible $(A,A)$-bimodules, $U\cong\operatorname{Hom}_{A}(V,A)$ and $V\cong\operatorname{Hom}(U,A)$ hold as isomorphisms of $(A,A)$-bimodules. Moving $AS^{-1}$ inside the parentheses is justified by [eisenbud, Proposition 2.10], the proof of which holds in the noncommutative setting. Recall that $V$ and $U$ are invertible bimodules, and are thus finitely generated and projective $A$ modules on the left and right. In particular, they are finitely presented as left and right $A$-modules.
(2)Notice that $V$ is finitely generated and projective as a left $A$-modules, so $S^{-1}V$ remains so as a left $S^{-1}A$-module. Now apply [sacred3, Section 2, Exercise 20].
@@
Let $A$ be a $k$-algebra, and let $S$ be a left and right denominator subset of $A$. If $A$ is twisted Calabi–Yau of dimension $d$, then $S^{-1}A$ is twisted Calabi–Yau of dimension $d$.
@@
$\mathfrak{S}^{-1}U\otimes_{S^{-1}A}\mathfrak{S}^{-1}V$ $\cong S^{-1}U\otimes_{S^{-1}A}S^{-1}V\cong S^{-1}(U\otimes_{A}V)\cong S^{-1}A.$
Recall that the first isomorphism is Lemma 4.6 applied to the same argument used in Lemma 4.3 to argue that $\mathfrak{S}^{-1}A\cong S^{-1}A$ as $(S^{-1}A)^{e}$-modules.